OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 1.18x + 4.34 = 8.67

OpenStudy (anonymous):

1.18x=8.67-4.34 1.18x=4.33 x=4.33/1.18 x=3.669492

OpenStudy (anonymous):

1.18x+4.34=8.67 -4.34 -4.34 1.18x=4.33 1.18 1.18 x=3.66 Answer is 3.7